# Evaluation of Dataframe Libraries for Data Preparation on a Single Machine

### v1

@@ IV-B Impact of Lazy Evaluation
Figure 8: Average runtime for running the entire pipeline on incremental samples of Taxi.
IV-C Scalability
@@
SummarySparkSQL is the best library for scaling to large datasets on a single machine. For both Patrol and Taxi datasets, SparkSQL is the only one able to execute the entire pipelines with the minimum configuration. CuDF follows closely but requires a GPU.
In this section, we assess how dataframe libraries scale by varying the size of the dataset and the configuration of the underlying machine. In particular, we use three Docker containers, whose CPU and RAM are incrementally increased as described in Table IV to reproduce typical configurations for laptops, workstations, and servers. Figure 8 show the performance of the libraries for running the entire pipeline on incremental samples of Taxi. Note that Pandas and Modin represent both their variants, since they show the same performance. Overall, SparkSQL stands out as the best performer and is the only one capable of handling the entire dataset on a laptop configuration (Figure 7(a)).
Moving to a workstation configuration (Figure 7(b)), even CuDF and Vaex can complete the pipeline, with the latter providing similar performance to SparkSQL. Finally, DataTable manages to complete the pipeline only on a server configuration (Figure 7(c)), while Pandas, SparkPD, Modin, and Polars all fail much earlier. In particular, Polars shows a good trend (similar to CuDF without the need for a GPU), but already faces memory issues to handle 15% of the dataset. Pandas and SparkPD exhibit the poorest performance instead, followed by Modin.
Table V outlines instead the minimum configuration required by each specific library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi. On both datasets, as previously stated, only SparkSQL can complete the entire pipeline with the minimum configuration. Following closely is CuDF, which can handle the entire Patrol dataset with a laptop configuration, but needs a GPU, making it more expensive. Further, Vaex stands out on Patrol as one of the libraries that require the least resources, followed closely by SparkPD. ModinR registers a performance on Taxi similar to its Dask counterpart, but scales much better on Patrol. In fact, while Dask engine uses a centralized scheduler to share work across multiple cores, Ray uses distributed bottom-up scheduling. Indeed, Modin can run on top of Dask, but it was originally built to work with Ray, and such integration is still more mature. DataTable maintains a relatively consistent performance on both datasets. Instead, despite excelling in other experiments, scalability appears to be a significant weakness for Polars. Finally, Pandas continues to show severe limitations.
@@
TABLE V: Minimum machine configuration for running the entire pipeline on incremental dataset samples. Laptop, Workstation, Server. See Table IV for the detailed configurations.
@@
ModinD LP LP WS WS WS SV SV LP LP WS WS SV
ModinR LP LP LP LP WS WS SV LP LP WS WS SV
Polars LP LP LP WS WS WS SV LP LP WS
cuDF LP LP LP LP LP LP LP LP LP LP WS WS WS WS
Vaex LP LP LP LP LP WS WS LP LP LP WS WS WS WS
Datatable LP LP LP WS SV SV SV LP LP LP WS WS WS SV
V Related Work
@@
To the best of our knowledge, our paper presents the first rigorous and extensive experimental comparison of existing dataframe libraries on data preparation tasks. In fact, public wisdom about such libraries is mostly scattered across several not peer-reviewed sources [16, 17, 18, 19, 20], which generally compare the performance of few libraries (e.g., Pandas vs PySpark vs Polars) on a handful of operations. The major claims by such analysis are mostly confirmed by our evaluation; nevertheless, despite often providing some useful insights, they only offer a partial and very fragmented knowledge, far from a complete and detailed overview covering all libraries and operations.
@@
In literature, papers presenting new dataframe libraries or dataframe-based tools usually include a comparison against some dataframe libraries. It is the case for the AFrame data analysis package [56] (later extended to PolyFrame [55]), designed to efficiently manage large-scale dataframes using SQL++ queries. AFrame is evaluated against Pandas, Spark, and Pandas on Ray (i.e., ModinR) through a micro-benchmark using the synthetic Wisconsin benchmark dataset [57] to assess acceleration and scalability across distributed environments. Similarly, Petersohn et al. [21] evaluate the benefits of Modin by comparing against Dask, Koalas, and Pandas. However, the mentioned comparisons share significant limitations, since data preparation is not the main topic of the analysis and only a relatively small set of operations is considered.
Further, dataframe libraries are sometimes compared in data science benchmarks. For instance, TPC-H [58], a popular decision support benchmark including a suite of business-oriented ad hoc queries and concurrent data modifications, is used by Polars [24] to compare against Pandas, PySpark, Dask, Vaex, and DuckDB [59]. Similarly, H2O (creator of DataTable[60] uses a set of queries to asses the scalability of group by and join operations on popular database and dataframe-like systems. Nevertheless, such relational benchmarking systems notoriously offer very limited insights about real-world workflows executed on dataframes [22].
@@
Sanzu [61] is instead designed to evaluate five popular data science frameworks and systems (R, Anaconda [62], Dask, PostgreSQL [63], and PySpark) on data processing and analysis tasks. Sanzu is composed of a micro-benchmark, to test basic operations in isolation on synthetic datasets of different sizes, and a macro-benchmark, to evaluate series of operations representing concrete data science scenarios on real-world datasets. FuzzyData [22] is a workflow generation system that allows to compare dataframe-based APIs on worflows composed of a small subset of operations. Finally, Shanbhag et al. [23] measure the energy consumption of 21 dataframe processing tasks on Pandas, Vaex, and Dask.
VI Key Takeaways
@@
Our evaluation confirmed the several known limitations of Pandas. However, despite the promises of the many libraries designed to overtake it, there is no silver bullet for data preparation of tabular data with dataframes on a single machine. We can help data practitioners to decide which library to use by answering the following questions.
VI-A Does the dataset fit in the GPU memory?
@@
If a GPU with enough memory to fit the data is available, CuDF appears to consistently yield the best overall performance and has quasi-complete compatibility with the Pandas API. However, CuDF lacks of an optimizer, thus libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations [20].
VI-B No complete pipeline execution?
@@
If a user is only interested in an individual stage (I/O, EDA, data transformation, or data cleaning), DataTable stands out as the fastest library for I/O operations—for other tasks it might not be that mature, with many preparators to implement manually. Polars emerges as the best performer for EDA. However, for Data Transformation and Data Cleaning, SparkSQL and Vaex show the best results.
VI-C What is the size of the datasets?
@@
With small datasets ($<500k$ rows), all of the libraries show similar average runtimes. Even if Polars is often recommended for such a scenario, Pandas (especially Pandas2) appears to be the most reasonable choice instead, since their performance does not differ substantially, at least not enough to learn a new API and replace entire parts of the pipeline [18]. In general, if memory and CPU usage are limited, Polars has the best overall performance.
@@
For datasets of medium size (from 2 to 19 million rows) with dense numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed. On the other hand, for sparse datasets with many null values (e.g., Loan), Polars and SparkPD show good performance. Finally, for datasets with many string columns (e.g., 25% Patrol), SparkSQL appears to be the best choice.
@@
For larger datasets (over 20 million rows), SparkSQL stands out as the best option [64]. Furthermore, when the pipeline requires dealing with heavy-duty column operations, particularly in datasets that have a multitude of string columns and a good amount of null values (e.g., Patrol), Vaex seems to be the best move—and there is no need to set up any environment, as required for PySpark.
VII Conclusion
We presented an experimental comparison of the most used dataframe-based libraries to help practitioners in selecting the best suitable to run their data preparation tasks on a single machine. To guarantee reasonably comparable results and increase usability, we developed Bento, a general framework for assessing library performance on four major data preparation stages: Input/Output (I/O), Exploratory Data Analysis (EDA), Data Transformation (DT), and Data Cleaning (DC). Our contribution includes a thorough analysis of performance of the most used dataframe libraries on heterogeneous datasets, that ranges in size, complexity, and characteristics and the relative most voted pipeline from Kaggle.com. The results of our experiments show that there is no clear winner among the libraries, thus we summarized useful takeaways for supporting the selection of the right library for a task and dataset at hand. As future work, we are planning to compare the libraries in a distributed setting and to study how machine learning can support users in suggesting optimal library combinations based on datasets, preparators, and previous executions.
Acknowledgment
@@
We extend our appreciation to the master’s students of the 2021/2022 academic year in the Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions to testing the libraries. This work was partially supported by the Department of Engineering “Enzo Ferrari” within the FARD-2022 (FAR2022DIP_DIEF - Prot. n. 2985 del 27/07/2022) and FARD-2023 (FAR_DIP_2023_DIEF-SIMONINI) projects.

### v2

@@ 4.2. Impact of Lazy Evaluation
Figure 6. Average runtime for running the entire pipeline on incremental samples of Taxi.
4.3. Scalability
{mdframed}
@@ 4.3. Scalability
SummarySparkSQL is the best library for scaling to large datasets on a single machine. Indeed, for both Patrol and Taxi, it is the only one able to execute the entire pipelines with the laptop configuration. Polars follows closely, but requires a lot of resources in terms of RAM, exhibiting the worst scalability on machines with limited resources.
In this section, we assess how dataframe libraries scale by varying the size of the dataset and the configuration of the underlying machine. In particular, we take into account three different machine configurations, whose CPU and RAM are incrementally increased as described in Table 4 to reproduce typical configurations for laptops, workstations, and servers. Figure 6 shows the performance of the libraries for running the entire pipeline on incremental samples of Taxi. Table 5 outlines instead the minimum configuration required by each library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi. Since CuDF scales according to GPU memory, it is not included in this experiment. If the dataset fits in the GPU, CuDF is able to efficiently handle it and complete the pipeline. However, this dependency does not make it a good candidate for dealing with very large datasets on machines with limited GPU resources.
Overall, SparkSQL clearly stands out as the best performer, being the only library capable of handling both datasets in their entirety on a laptop configuration. This capability is due to the use of disk spillover, which offloads data from RAM to disk storage, when the memory limit is reached during computations, ensuring continuous operation without exceeding memory capacity. On the other hand, SparkPD suffers instead from overhead due to internal conversions between Pandas and Spark dataframes and increased JVM memory usage, causing out-of-memory errors on the laptop configuration for both Patrol and Taxi.
DataTable stands out on Patrol as one of the libraries that require the least resources, followed closely by ModinR and Pandas. Indeed, it is one of the libraries that consumes the least memory on Patrol; further, it does not require to split the dataframe into multiple pieces, which can lead to errors due to communication and information exchange.
ModinR demonstrates better scalability across all machine configurations compared to its Dask counterpart. The Dask engine uses a centralized scheduler to distribute data across multiple cores, which results in higher memory consumption and leads to out-of-memory (OOM) issues more easily. In contrast, Ray can scale far beyond Dask due to its distributed task scheduling scheme, specifically employing a distributed bottom-up scheduling approach. This optimizes resource utilization and minimizes overhead by initiating tasks at the lowest level of computation and aggregating results upwards. Notably, Modin was originally built to work on top of Ray, making this integration more mature and optimized.
Despite its excellent performance in other experiments, scalability appears to be a significant weakness for Polars. This limitation arises from its in-memory execution model, which requires all data to be loaded into RAM. While this approach offers speed and efficiency for smaller datasets, it becomes a bottleneck when handling larger datasets, limiting its scalability.
Finally, Pandas confirms its well-known severe limitations about scalability, being the only library not able to complete the pipeline on Taxi even on a server configuration (Figure 6(c)).
@@
Table 5. Minimum machine configuration for running the entire pipeline on incremental dataset samples. Laptop, Workstation, Server. See Table 4 for the detailed configurations.
@@
ModinD LP LP LP WS SV SV SV LP LP LP WS SV
ModinR LP LP LP LP WS WS WS LP LP LP LP WS WS WS
Polars LP LP LP WS WS WS WS LP LP WS WS WS SV SV
Vaex LP LP LP WS WS WS WS LP LP LP WS WS WS WS
Datatable LP LP LP LP LP WS WS LP LP LP LP WS WS WS
4.4. Performance on the TPC-H queries
{mdframed}
SummaryCuDF consistently achieves the best performance across all queries, while Polars significantly outperforms other CPU-only dataframe libraries.
TPC-H [^27] https://www.tpc.org/tpch/default5.asp is a decision support benchmark consisting of a suite of business-oriented ad-hoc queries and concurrent data modifications (Pöss and Floyd, 2000). It is widely adopted to compare end-to-end database systems (Dreseler et al., 2020), allowing to comprehensively evaluate their performance and efficiency through the execution of complex queries on large volumes of data. Consistently with previous related work [^28] https://pola.rs/posts/benchmarks/ exploiting TPC-H queries to assess the performance of a subset of dataframe libraries (see Section 5 for more details), we also adopt this benchmark to provide further support to the outcomes of the comparative analysis presented above.
In particular, we take as a reference a publicly available translation of the TPC-H queries into Pandas API operations [^29] https://github.com/Bodo-inc/Bodo-examples, replicating it for the other libraries, then use Bento for their seamless execution. Figure 7 illustrates the performance of each library on the 22 queries of the TPC-H 10GB benchmark. We select a scale factor of 10, because it represents the largest dataset that can be processed by the 40 GB RAM of our GPU, which therefore would not be able to handle a scale factor of 100.
Overall, CuDF consistently achieves the best performance by leveraging GPU power. Among the CPU-only libraries, Polars is instead the clear winner. SparkSQL significantly outperforms Pandas on several queries, registering notable results especially on q06, where it is even faster than Polars (note that this query selects line items shipped within a year interval, reducing the dataset by 85%). Pandas is not the worst performer, but on most queries it does not achieve good results. Modin (which registers the same results with both engines, Dask and Ray) shows diverse performance instead, strongly dependent on the query.
As mentioned above, SparkPD suffers from the additional latency due to the translation of Pandas operations into Spark execution plans, making it one of the worst performers. DataTable is one of the worst performers too, since it is slow in grouping operations (as also testified by their own benchmark [^30] https://h2oai.github.io/db-benchmark/) and its API only supports joins on columns with unique values, requiring therefore to switch to the default Pandas API (e.g., q09). Finally, Vaex is by far the worst performer overall, as it is significantly slow in grouping operations (as also pointed out in the previous sections) and lacks support for multi-column joins.
5. Related Work
@@
To the best of our knowledge, our paper presents the first rigorous and extensive experimental comparison of existing dataframe libraries on data preparation tasks. In fact, public wisdom about such libraries is mostly scattered across several not peer-reviewed sources (Schmitt, 2020; Alexander, 2023; Pinner, 2023), which generally compare the performance of few libraries (e.g., Pandas vs PySpark vs Polars) on a handful of operations. The major claims by such analysis are mostly confirmed by our evaluation; nevertheless, despite often providing some useful insights, they only offer a partial and very fragmented knowledge, far from a complete and detailed overview covering all libraries and operations.
@@
In literature, when a paper introduces a new dataframe library or dataframe-based tool, it usually includes a comparison against some dataframe libraries. It is the case for the AFrame data analysis package (Sinthong and Carey, 2019) (later extended to PolyFrame (Sinthong and Carey, 2021)), designed to efficiently manage large-scale dataframes using SQL++ queries. AFrame is evaluated against Pandas, Spark, and Pandas on Ray (i.e., ModinR) through a micro-benchmark using the synthetic Wisconsin benchmark dataset (DeWitt, 1993) to assess acceleration and scalability across distributed environments. Similarly, Petersohn et al. (Petersohn et al., 2021) evaluate the benefits of Modin by comparing against Dask, Koalas, and Pandas. However, these comparisons share significant limitations, since data preparation is not their main focus and only a relatively small set of libraries and operations is considered.
Further, dataframe libraries are sometimes compared in data science benchmarks. For instance, as previously mentioned in Section 4.4, it is the case for TPC-H (Pöss and Floyd, 2000), which is used by Polars  [^31] https://pola.rs/posts/benchmarks/ to compare against Pandas, PySpark, Dask, Modin, and DuckDB (Raasveldt and Mühleisen, 2019). While the outcome of their analysis is consistent with our findings, their evaluation (still performed by selecting a scale factor of 10) only covers the first 7 queries out of the entire batch of 22 covered by our evaluation. Similarly, H2O (creator of DataTable) uses a set of queries to asses the scalability of popular database and dataframe-like systems [^32] https://h2oai.github.io/db-benchmark only limited to group by and join operations. Indeed, due to their different focus, such relational benchmarking systems notoriously offer very limited insights into real-world data preparation workflows executed on dataframes (Rehman and Elmore, 2022). Note that for producing a consistent a analysis we do not take into account tools that are not based on the dataframe data structure, as the aforementioned DuckDB or similar SQL-based solutions. For instance, Grizzly (Kläbe and Hagedorn, 2021) aims to address Pandas limitations in scalability by translating dataframes to SQL code, while PyFroid translates from the Pandas API to the DuckDB API to efficiently scale Pandas workloads on a commodity workstation, using top-voted Kaggle notebooks for the evaluation.
@@
Sanzu (Watson et al., 2017) is designed to evaluate five popular data science frameworks and systems (R, Anaconda [^33] https://www.anaconda.com, Dask, PostgreSQL [^34] https://www.postgresql.org, and PySpark) on data processing and analysis tasks. Sanzu is composed of a micro-benchmark, to test basic operations in isolation on synthetic datasets of different sizes, and a macro-benchmark, to evaluate series of operations representing concrete data science scenarios on real-world datasets. FuzzyData (Rehman and Elmore, 2022) is a workflow generation system that allows to compare dataframe-based APIs on workflows composed of a small subset of operations. Finally, Shanbhag et al. (Shanbhag and Chimalakonda, 2023) measure the energy consumption of 21 dataframe processing tasks on Pandas, Vaex, and Dask.
(a) Queries 1 to 11.
(b) Queries 12 to 22.
Figure 7. Performance of the dataframe libraries on the TPC-H 10GB queries.
6. Key Takeaways
@@
Our evaluation confirmed the well-known limitations of Pandas. However, despite the abundance of alternative dataframe libraries designed to overtake it, there is no silver bullet to perform data preparation of tabular data with dataframes on a single machine. The right choice usually depends on many factors, such as the size and the features of the dataset at hand, the configuration of the underlying machine, and the kind of operations to carry out. We can help data practitioners to decide which library to use by answering the following questions.
Does the dataset fit in the GPU memory?
@@
If a GPU with enough memory to fit the data is available, CuDF appears to consistently yield the best overall performance and has quasi-complete compatibility with the Pandas API. However, CuDF lacks of an optimizer; thus, libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations [^35] https://pola-rs.github.io/polars/user-guide/misc/alternatives/.
Individual stage or complete pipeline?
@@
If a user is only interested in an individual stage, i.e., input/output (I/O), explorative data analysis (EDA), data transformation (DT), or data cleaning (DC), DataTable stands out as the fastest library for I/O operations, while for other tasks it might not be that mature. Further, it shows the least compatibility with Pandas, requiring to implement many preparators manually. Polars emerges as the best performer for EDA, while SparkSQL represents a consistent solution for DT and DC, also recording the best performance among all CPU-only libraries when running the entire pipeline.
What is the size of the datasets?
@@
For small datasets (less than 500k rows), all of the libraries show similar average runtimes. Even if Polars is often recommended for such a scenario, and indeed registers the best overall performance if memory resources are not limited, Pandas might still be the most reasonable choice, since their performance does not differ substantially—at least, not enough to learn a new API and replace entire parts of the pipeline (Pinner, 2023).
@@
In the case of datasets of medium size (from 2 to 19 million rows) with dense numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed. On the other hand, for sparse datasets with many null values (e.g., Loan), Polars and SparkPD show good performance. Finally, for datasets with many string columns (e.g., 25% Patrol), SparkSQL appears to be the best choice.
@@
For larger datasets (over 20 million rows), SparkSQL stands out as the best option (Tang, 2023). Furthermore, when the pipeline requires dealing with heavy-duty column operations, particularly in datasets that have a multitude of string columns and a good amount of null values (e.g., Patrol), Vaex seems to be the best move—and there is no need to set up any environment, as required for PySpark.
7. Conclusion
We presented a comprehensive experimental comparison of the most used dataframe libraries to support practitioners in the selection of the most suitable solution to carry out their data preparation tasks on a single machine. To guarantee comparable results and increase usability, we developed Bento, a general framework for assessing the performance of dataframe libraries on four major data preparation stages: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC). We exploited Bento to perform a thorough comparative analysis of the libraries on heterogeneous datasets (varying in size, complexity, and features) previously adopted in literature and publicly available on Kaggle. In particular, we relied on the most voted Kaggle notebooks to assess the performance of the libraries on real-world data preparations pipelines validated by a large community. Further, we also compare their performance on the popular TPC-H benchmark to support our findings.
The results of our experiments show that there is no clear winner among the libraries, whose performance is strongly dependent on the size and features of the dataset at hand, on the specifications of the underlying machine, and on the kind of operations to carry out. Thus, we provided the readers with useful takeaways to support them in the selection of the right library for the data preparation task at hand. As future work, we plan to compare the libraries in a distributed environment and study the possible adoption of machine learning to suggest optimal library combinations based on datasets, tasks, hardware, and previous executions.
Acknowledgements
@@
We extend our appreciation to the master’s students of the 2021/2022 Big Data Management and Governance course at the University of Modena and Reggio Emilia for their contributions in testing the libraries. We warmly thank Leonardo S.p.A., co-funding the PhD program of Angelo Mozzillo. Further, this work was partially supported by the “Enzo Ferrari” Engineering Department (University of Modena and Reggio Emilia) within the projects FARD-2022 and FARD-2023, and by MUR within the project “Discount Quality for Responsible Data Science: Human-in-the-Loop for Quality Data” (code 202248FWFS).